OpenStudy (anonymous):

For the function f(x) = 3x + 3, find x such that f(x) = 18 A. 6 B. 57 C. 5 D. 3

geerky42 (geerky42):

So you have \(18= f(x) = 3x+3 \rightarrow18=3x+3\) Can you solve \(18=3x+3\)?